Compound 2: Use the information provided below and the IR and NMR spectra on the next page to answer the following questions. a) An MS was taken of compound 2 and the table is given below. Determine the molecular formula of 2 from the MS given below relative abundance m/z 132 100 133 9.8 b) Calculate the Index of Hydrogen Deficiency/Degree of Unsaturation for the molecule. c) Using the IR spectrum and the 1H- and 13C-NMR spectra given below, determine a structure of this unknown molecule that is consistent with all data. Draw the structure of compound 2 in the box below.
wVEHP S 1H 2H 3H 1H d 1H 10 PPM 3 signals here 160 100 PPM 120 60 200 40 180 140 80 20 E un-
Answer #1

a) malecular formula: CaHeo mumber of Carbong =9- Hs=122-(12x9) - (ixi) = 8 by Degree of unlafuraton 20 +2 +N- X-H 2 2 (9)+24-NMR .CH=CH-CH 89.5 86-7.d H m d,2H 8구6 873, 3H IR ~3030 Ch-Arc-H Str 1690 cn C=D Str C-H St 2구00, 2600 Cm